what is algebric function

Grade:11

1 Answers

Saurabh Singh
10 years ago

Inmathematics, analgebraic functionis afunctionthat can be defined as theroot of a polynomial equation. Quite often algebraic functions can be expressed using a finite number of terms, involving only thealgebraic operationsaddition, subtraction, multiplication, division, and raising to a fractional power:

f(x)=1/x, f(x)=\sqrt{x}, f(x)=\frac{ \sqrt{1+x^3}}{x^{3/7}-\sqrt{7} x^{1/3}}

are typical examples.

Thanks & Regards

Saurabh Singh,

askIITians Faculty

B.Tech.

IIT Kanpur